Find the Integration of limit (0 to 1) 1/(1+x^8) .dx

Grade:12th pass

3 Answers

Khushan Sanghvi
5 years ago
∫1/1+(x^4)^2= arctanx^4  when u apply limits u get pi/4 as final ans..!!!
Sai Ram Charan
5 years ago
Hello Kaushik! did you mean the following integral?
\int_{0}^{1}\frac{1}{1+x^8}dx
 this integral doesnot appear in class 12 syllabus at all!
Sai Ram Charan
5 years ago
But, if you still need an answer, i would give a solution as:
\left(\displaystyle\sum_{\left\{Z:\>Z^8+1=0\right\}}\dfrac{\ln\left(\left|x-Z\right|\right)}{8Z^7}\right)   
I could not simplify it further. Sorry!